When the reciprocal of three times a number is subtracted from the reciprocal of the number, the result is one-sixth. Find the number.
Found 2 solutions by richwmiller, edjones:
Answer by richwmiller(17219)

reciprocal means 1 over the number
a number times its reciprocal =1
x*1/x=1
the reciprocal of x=1/x
1/x-1/3x=1/6

Answer by edjones(8007)

Let x be the number.
1/x - 1/3x = 1/6
6-2=x
x=4
.
Ed